# Krylov Complexity in early universe

### v2

@@ 5.2 Krylov Entropy
where $\phi_{n}$ is the wave function (27). Once given the analytic formula of K-entropy, we could numerically simulate it in various epochs as shown in Fig. 5.
@@
Fig. 5 indicates a complete evolution of K-entropy in inflation, RD and MD, which is similar to Krylov complexity. We could see that the curvature of perturbation will be more and more disorder in inflation. For RD and MD, the disorder will be enhancing and then saturates to some constant values and the total trend is enhancing, where the trend is also highly related the scale factor. From another aspect, the preheating will generate the particles occurred at MD that leads to the growth of K-entropy.
@@
Let’s discuss Section 5. In this section, we provide a comprehensive analysis of Krylov complexity throughout the early universe. Our numerical data clearly indicates that different potentials yield almost identical patterns of Krylov complexity and K-entropy. Specifically, Krylov complexity increases during inflation and then stabilizes at constant values during RD and MD. This same behavior is observed for K-entropy. Additionally, we must examine the Lanczos coefficient (equivalent to the Lyapunov index in maximal chaos), which remains consistent in both the closed and open system methods.
6 Krylov complexity with the approach of open system
@@
The preheating process in MD and RD is a highly non-equilibrium process that creates particles Kofman:1997yn ; Shtanov:1994ce . In addition, Ref. Cheung:2007st demonstrated that the universe’s evolution breaks the time invariance when the temporal direction is reversed, leading to a violation of energy conservation. Due to these reasons, the open system approach is considered to be more realistic compared to the closed system method. We will follow the method of Bhattacharya:2022gbz to investigate the Krylov complexity.
6.1 Lanczos coefficient and dissipation coefficient
Firstly, the general operator in Heisenberg’s picture has the following form,
$\mathcal{O}(\eta)=e^{i\mathcal{L}_{o}\eta}$ (42)
where the Lindbladian $\mathcal{L}_{o}$, also known as the Hamiltonian, acts on the Krylov basis, we can obtain
$\mathcal{L}_{o}|\mathcal{O}_{n})=-ic_{n}|\mathcal{O}_{n})+b_{n+1}|\mathcal{O}_{n+1})+b_{n}|\mathcal{O}_{n-1})$ (43)
where $c_{n}$ encodes the information of the open system and $b_{n}$ is the Lanczos coefficient. Due to this extra term, we could divide the total Hamiltonian (also known as Lindbladian) into two parts which can be represented by
@@ 6.1 Lanczos coefficient and dissipation coefficient
In our work, $\mathcal{L}_{o}$ is the Hamiltonian Eq. (33). In light of (43), we can easily obtain the open part (generating $c_{n}$) of Hamiltonian,
@@
where $|\mathcal{O}_{n})=\Bigl{|}n_{\vec{k}};n_{-\vec{k}}\Bigr{\rangle}$. By substituting Eq. (45) and Eq. (46) into Eq. (43), we can obtain the equations for $c_{n}$ and $b_{n}$ as
@@
Here, we will discuss more about $b_{n}$ and $c_{n}$. Ref. Bhattacharjee:2022lzy has introduced the $a_{n}$ corresponding to $c_{n}$ in our work and also stated that $c_{n}$ is highly related to the dissipative coefficient. Meanwhile, we could see that $b_{n}$ is the same for the closed system and open system since it is precisely determined by the Hamiltonian.
In Ref. Parker:2018yvk , it has clearly indicated that the dynamical system is an infinite, non-integrable, many-body system as $b_{n}=\alpha n+\gamma$, where $\alpha$ and $\gamma$ are two parameters determined by various models. In our case, we could see that $b_{n}\propto n$, thus the whole early universe is an ideal chaotic system.
6.1.1 $b_{n}$ and $\mu_{2}$
[…]
@@ 6.1.1 $b_{n}$ and $\mu_{2}$
where we have defined $\tilde{c}_{n}=-ic_{n}$, meanwhile we have used the $b_{n+1}\approx b_{n}$ and $\phi_{n+1}-\phi_{n-1}\approx 2\partial_{n}\phi$ in the continous limit. $b_{n}$ and $c_{n}$ are defined in Eq. (47) and Eq. (48), respectively. The generic formulas of $b_{n}$ and $c_{n}$ are
@@
where $\mu_{1}$, $\mu_{2}$ and $\beta$ are determined by various models and $\mu_{2}$ plays a role of dissipative coefficient as shown in Ref. Li:2024kfm . In our work, one can explicitly find $\beta=1$ and combine Eqs. (47), (48), and (50), we find that the parameters $\mu_{1}$ and $\mu_{2}$ have the following forms,
@@
It should be noticed that $\mu_{2}$ is mainly determined by $k$ in inflation due to the slow-roll conditions. Once entering RD and MD (taking the contribution of the effective mass into account), we could see that it will be smaller than unity since the effective mass of these three types potentials is much more tiny compared with momentum $k$. First, we discuss the Lanczos coefficient $b_{n}$. Fig. 6 clearly indicates the three distinctive potentials of $b_{n}$ in inflation, RD and MD, in which the parameters of $k$, $H_{0}$ are set to be the same with the previous plots. As Bhattacharya:2022gbz confirmed, the Lyapunov index is equavillent to $b_{n}$ whose relation is $\lambda=2\alpha$ with $b_{n}=\alpha n+\delta$ ($\delta=0$ in our case), where it is explicitly showing that $\lambda=2\alpha$, namely $\lambda=b_{n}$ with $n=2$ which will be shown in Fig. 6. The physical meaning of $\lambda$ represents the chaotic feature of dynamical system. Consequently, $b_{n}$ in our case also manifest the chaotic feature for the early universe.
@@
Figure 6 illustrates the evolution trend $b_{n}$ in inflation, radiation domination (RD), and matter domination (MD). Based on the discussion regarding $b_{n}$, chaos will be heightened due to the exponential expansion of the background space. As the energy decreases, $b_{n}$ will also decline since the temperature decreases. During MD, particles are generated through a preheating process known as parametric resonance, which leads to an increase in $b_{n}$. Finally, we need to analyze the dissipative coefficient $\mu_{2}$ that is related to $c_{n}$, where we could confirm the previous results via Li:2024kfm . Fig. 7 clearly shows the evolution in RD and MD. The $\mu_{2}$ in inflation is not shown in Fig. 7 since $\mu_{2}=k=1$ in light of the values of $k$ for inflation. Ref. Li:2024kfm claimed that inflation is a strong dissipative system as $\mu_{2}\geq 1$, RD and MD are weak dissipative systems as $\mu_{2}\ll 1$, which is explicitly confirmed by Fig. 7.
@@
6.2 General discussion of $\phi_{n}$
Before the Krylov complexity, we discuss the general properties of $\phi_{n}$ since the definition of Krylov complexity is explicitly represented by $C_{K}=\sum_{n=1}^{+\infty}n|\phi_{n}|^{2}$. We will follow the general discussion of $\phi_{n}$ via Refs. Li:2024kfm ; Bhattacharjee:2022lzy . Eq. (49) can be rewritten by
@@ 6.2 General discussion of $\phi_{n}$
where we have defined $\tilde{c}_{n}=n\chi\mu$ and $b_{n}=n\alpha$. As $\eta$ approaches infinity, the solution of Eq. (52) has the form $\phi(n)\propto e^{-n\xi}$ with $\xi=\frac{2\alpha}{\mu\chi}$, where this form will appear in RD or MD due to their weak dissipative nature. In the weak dissipative region, the corresponding Krylov complexity behaves as follows: it increases exponentially at the onset, with the average position given by $C_{k}\propto e^{2\alpha\eta}$ before reaching $\eta_{*}$. After this point, the Krylov complexity saturates to a constant value $\xi$ when $\eta>\eta_{*}$, where $\eta_{*}\sim\frac{1}{2\alpha}\ln{\frac{2\alpha}{\chi\mu}}$ dubbed as the scrambling time. It is evident that $\eta_{*}$ determines the range of exponential growth of Krylov complexity, determined by the dynamical system itself. It is straightforward to demonstrate that $\exp(2a\eta_{*})=\xi$. After some simple algebra, we could obtain the formula $\xi$ in our case as follows,
@@
From this formula, we can observe that $\xi$ is not strictly constant; it varies with respect to $a$ and $k$. Consequently, a significant change in the background will have a substantial impact on Krylov complexity. In this section, we examined the general properties of Krylov complexity without providing the precise formula for $\phi_{n}$. To fully capture the information concerning Krylov complexity within the framework of the open system method, it is essential to have the exact wave function for $\phi_{n}$.
6.3 Wave function within the method of open system
Although we constructed the wave function for an open system in reference Li:2024kfm , based on the framework established in Bhattacharjee:2022lzy , our approach is too simplistic to capture the complete information of the open system. Notably, our wave function, as shown in Li:2024kfm , can effectively reduce to the case of a closed system, establishing a correspondence between $\eta$ and $r_{k}$. To clearly illustrate this issue, we will first present our constructed wave function as follows:
@@ 6.3 Wave function within the method of open system
To perform the Taylor expansion around $\mu_{2}$ to the first order, we obtain the expression
$\phi_{n}\approx|1-\mu_{1}^{2}|^{n/2}\,\text{sech}(\eta)\,\tanh^{n}(\eta)+\mathcal{O}(\mu_{2}^{n}).$
Next, we can neglect the term involving $\mu_{1}^{2}$, which enables us to find that this approximation matches the wave function given in equation (39) for $\phi_{k}=0$. Additionally, we observe the correspondence $\eta\sim r_{k}$. In principle, the wave function $\phi_{n}$ should incorporate the information of both $r_{k}$ and $\phi_{k}$ to accurately reproduce the wave function represented in (39) as the leading order of the open system’s method. To achieve this, we will follow the guidelines provided in Appendix D of Parker:2018yvk .
@@
where $x$ represents the Hamiltonian, $\tilde{c}_{n}$ is associated with the dissipative coefficient, and $b_{n}$ refers to the Lanczos coefficient. The sequence begins with the initial conditions $P_{0}(x)=1$ and $P_{1}(x)=x-c_{0}$. For an open system, we have $P_{n}(x;\alpha,\eta)=\det(x-\mathcal{L}_{n})$, where $\mathcal{L}_{n}$ is the Hamiltonian for the $n$-th quantum state. This determinant encodes information about $b_{n}$ as represented by the corresponding matrix. Based on the definition of $L$, we can express it in terms of $x^{n}$ or $e_{n}$, which will be used to derive the wave function using the open system method.
To obtain the similar recurrence relation as (43), we introduce the the natural orthonormal basis $\{e_{n}\}$, and meanwhile $P_{n}$ can be represented in terms of $e_{n}$
@@
where $x$ is the parameter of $P_{n}$, $\mathcal{L}$ represents our Hamiltonian expressed in terms of creation and annihilation operators, and $|e_{n}\rangle\equiv|\mathcal{O}_{n}\rangle=\Bigl{|}n_{\vec{k}};n_{-\vec{k}}\Bigr{\rangle}$. By combining Equation (55) with Equation (56), we can derive the following recurrence relation.
@@
where $\mu_{n,k}$ refers to a triangular linear transform with matrix elements that relate $x$ to $P_{n}$. This formula demonstrates that the polynomials $P_{n}$ can be expressed in terms of $x$. All three orthogonal polynomial sequences (OPS) mentioned are examples of Sheffer orthogonal polynomials, which are defined as a set of polynomials $\{P_{n}(x)\}_{n\geq 0}$ characterized by a generating function,
@@
In our calculations, we multiply a phase factor $ie^{2i\phi_{k}(\eta)}$ which does not alter the wave function. Finally, we can set $r_{k}=\eta$, allowing us to denote the two-mode squeezed state within the open system as,
@@
This wave function is our central result in this paper whose leading order of $\phi_{n}$ in terms of $\mu_{2}$ is consistent with the wave function described in (39). The idea of changing the variable is borrowed from the construction of coherent states, where a key component is the generation of the displacement operator, defined as $D(\eta)=e^{iHt}$, with $H$ being the corresponding Hamiltonian, where one can change $t$ in terms of $r_{k}$ and $\phi_{k}$. In Ref. Caputa:2021sib , it is stated that the generalized coherent state is derived by constructing a generalized displacement operator according to their group representation theory.
6.4 The power of wave function (73)
We discuss the two-mode squeezed state in an open system, as described in equation (73). Unlike the two-mode squeezed state derived from closed system methods, our wave function (73) is more realistic, as it considers a vast open environment. Additionally, although our derivation pertains to the early universe, this wave function can also be applied to fields such as quantum optics, quantum information, and condensed matter physics, where information is encoded in the parameters $\mu_{1}$ and $\mu_{2}$ determined by the specific Hamiltonian. Finally, we observe that our calculations hold true by requiring that the recurrence relation (55) which is equivalent to the Lanczos algorithm. This connection highlights the power of the Lanczos algorithm, which provides a natural method for constructing the wave function according to their Hamiltonian, even for open systems.
6.5 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$ in open system’s method
According to the standard procedure, we need to derive the evolution equations for $r_{k}$ and $\phi_{k}$ in relation to conformal time $\eta$. The detailed calculations for this derivation can be found in Appendix A. The evolution can be summarized as follows:

### v3

@@ V.2 Krylov Entropy
where $\phi_{n}$ is the wave function (28). Once given the analytic formula of K-entropy, we numerically simulate it in various epochs as shown in Fig. 5.
@@
Fig. 5 illustrates the complete evolutionary profile of the K-entropy across the inflationary, RD, and MD eras, exhibiting a trend similar to that of the Krylov complexity. It can be observed that the curvature perturbations become progressively more disordered during inflation. In the subsequent RD and MD phases, the degree of disorder initially increases before saturating to approximately constant values; nonetheless, the overall trend remains one of growth, which is closely tied to the evolution of the scale factor. From another perspective, preheating, which generates particles during the MD era, also contributes to the observed increase in the K-entropy.
@@
We now summarize the findings of Section V. In this section, we have conducted a comprehensive analysis of Krylov complexity in the early universe. Our numerical results clearly demonstrate that different inflationary potentials produce nearly identical patterns of evolution for both the Krylov complexity and the K-entropy. Specifically, the Krylov complexity grows during inflation and subsequently saturates to nearly constant values throughout the RD and MD eras. The K-entropy exhibits the same qualitative behavior. To characterize the chaotic features of the early universe more precisely, we must examine the Lanczos coefficients, which, in maximal chaos, play a role analogous to the Lyapunov exponent. These coefficients are identical in form for both the closed- and open-system approaches. In the following, we will employ the open-system methodology to investigate Krylov complexity.
VI Krylov complexity with open-system methodology
@@
The preheating process during the MD and RD eras is a highly non-equilibrium phenomenon that generates particles Kofman:1997yn ; Shtanov:1994ce . Furthermore, Ref. Cheung:2007st demonstrated that the evolution of the universe breaks time-reversal invariance, leading to a violation of energy conservation. For these reasons, the open-system methodology is regarded as more realistic than the closed-system approach. Several open-system methodologies exist for investigating Krylov complexity. Here, we will implement the generalization of the Lanczos algorithm to the Arnoldi iteration via the Heissenberg form Bhattacharjee:2022lzy ; Bhattacharya:2022gbz ; Liu:2022god ; Nizami:2023dkf ; Nandy:2024evd . Let us review this Arnoldi iteration step by step. We begin by recalling the general Lindblad master equation Lindblad:1975ef :
$\dot{\rho}=-i[H,\rho]+\sum_{k}[L_{k}\rho L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\rho\}],$ (44)
where $\rho$ is the density matrix, $H$ is the Hamiltonian, and $L_{k}^{\dagger}$ denotes the jump operator (which in our context corresponds to the annihilation operator). Following [81], we consider the open-system dynamics in an infinite-temperature environment. Specifically, we identify the system state $\rho$ with the maximally mixed state (infinite temperature state) $\rho_{\infty}$ due to the extremal high temperature for the early universe. This assumption defines the inner product for the operator space as $(A,B)=\text{Tr}(\rho_{\infty}A^{\dagger}B)$, which is central to our Krylov complexity calculations. Under this approximation, we obtain
$\sum_{k}[L_{k}\rho_{\infty}L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\rho_{\infty}\}]=0,$ (45)
In this stationary state, it is reasonable to assume that the operator dynamics is governed by
$\mathcal{\hat{O}}(t)=\exp(i\mathcal{L}t)\mathcal{\hat{O}}$, (46)
where the Heisenberg-picture Lindbladian $\mathcal{L}$ is given by Bhattacharya:2022gbz ,
$\mathcal{L}=\mathcal{L}_{H}+\mathcal{L}_{D},~~\mathcal{L}_{H}\mathcal{\hat{O}}=[H,\mathcal{\hat{O}}],~~\mathcal{L}_{D}=\sum_{k}[L_{k}\mathcal{\hat{O}}L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\mathcal{\hat{O}}\}$] (47)
where $\mathcal{L}_{H}$ governs the Hermitian dynamics, $\mathcal{L}_{D}$ denotes the dissipative part, and $\mathcal{L}$ behaves as a superoperator on the operator space. The inner product is defined as $(A|B):=\mathrm{Tr}[\rho_{\infty}A^{\dagger}B]$, which explicitly connects the operator space structure to the Krylov basis,
$\rm span(\mathcal{O}_{0},...,\mathcal{O}_{n})=span(\mathcal{V}_{0},...,\mathcal{V}_{n}),$ (48)
where $\mathcal{V}_{n}$ denotes the original Krylov basis, which is generated via the Arnoldi iteration. In this work, we follow the standard construction, beginning with an initial normalized operator $\mathcal{O}_{0}\propto\mathcal{O}$. Subsequent Krylov basis vectors are then generated by repeatedly applying the Liouvillian superoperator and performing orthonormalization at each step. To compute the physical quantity, an appropriate basis is required. Following Bhattacharya:2022gbz , we adopt an orthonormal basis $\{\mathcal{O}_{0},\ldots,\mathcal{O}_{n},\ldots\}$ generated by the open-system Lindbladian:
$\rm span(\mathcal{O}_{0},...,\mathcal{O}_{n})=span(\mathcal{V}_{0},...,\mathcal{V}_{n}),$ (49)
where $\mathcal{V}_{n}$ is the original Krylov basis, and we generate this basis by using the Arnoldi iteration, which is started with an initial normalized vector $\mathcal{O}_{0}\propto\mathcal{O}$. Thereafter, we construct the subsequent Krylov basis vectors by repeating the following steps:
$|\mathcal{U}_{k})=\mathcal{L}|\mathcal{O}_{k-1}).$ (50)
From $j=0$ to n1n-1, we perfom the following iterations: $1.~h_{j,k-1}=(\mathcal{O}_{j}|\mathcal{U}_{k})$, $2.~|\tilde{\mathcal{U}}_{k})=|\mathcal{U}_{k})-\sum_{j=0}^{k-1}h_{j,k-1}|\mathcal{O}_{j})$, $3.~h_{k,k-1}=\sqrt{(\tilde{U}_{k}|\tilde{\mathcal{U}}_{k})}$, it will stop as if $h_{k,k-1}=0$, otherwise, it is defined as follows,
$|\mathcal{O}_{k})=\frac{|\tilde{\mathcal{U}}_{k})}{h_{k,k-1}}.$ (51)
Finally, the Lindbladian can transform into an upper Heissenberg form in the span Krylov basis:
$\mathcal{L}=\begin{pmatrix}h_{0,0}&h_{0,1}&h_{0,2}&\cdots&\cdots&h_{0,n}\\ h_{1,0}&h_{1,1}&h_{1,2}&\cdots&\cdots&h_{1,n}\\ 0&h_{2,1}&h_{2,2}&h_{2,3}&\cdots&\cdots\\ \vdots&0&h_{3,2}&\cdots&\cdots&\cdots\\ 0&\vdots&0&\vdots&\ddots&h_{n-1,n}\\ 0&0&\vdots&0&h_{n,n-1}&h_{n,n}\end{pmatrix}$ (52)
where $h_{m,n}=(\mathcal{O}_{m}|\mathcal{L}|\mathcal{O}_{n})$. As $\mathcal{L}$ is a Hermitian superoperator, the Arnoldi iteration will reduce into the Lanczos algorithm. Following conventions of Li:2024kfm , we represent the generalized Lanczos algorithm as
$\mathcal{L}|\mathcal{O}_{n})=-ic_{n}|\mathcal{O}_{n})+b_{n+1}|\mathcal{O}_{n+1})+b_{n}|\mathcal{O}_{n-1}),$ (53)
where the coefficient $c_{n}$ encodes the information of the open quantum system dynamics for a given Hamiltonian, corresponding to the diagonal elements $h_{n,n}$ in Eq. (52). Meanwhile, the Lanczos coefficients $b_{n}$, associated with the off-diagonal elements $h_{n,n-1}$ in Eq. (52), characterize the chaotic behavior of the dynamical system. The orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$ spans the Krylov space defined by Eq. (49). The Liouvillian superoperator $\mathcal{L}$ is equivalent to a Hamiltonian operator that admits a representation in terms of creation and annihilation operators.
VI.1 Lanczos coefficient and dissipation coefficient
Based on the foregoing analysis, we separate the total Hamiltonian (the Lindbladian) into two constituent parts:
@@ VI.1 Lanczos coefficient and dissipation coefficient
In light of (53), we easily obtain the open part (generating $c_{n}$) of Hamiltonian,
@@
where $|\mathcal{O}_{n})=\Bigl|n_{\vec{k}};n_{-\vec{k}}\Bigr\rangle$. Specifically, the jump operators $L_{k}$ appearing in the dissipative term $L_{D}$ are identified with the annihilation and creation operators in our work, i.e., $L_{k}=c_{k}$ and $L_{-k}=c_{-k}^{\dagger}$. By substituting Eq. (55) and Eq. (56) into Eq. (53), we obtain the equations for $c_{n}$ and $b_{n}$ as
@@
Here, we provide further discussion regarding the coefficients $b_{n}$ and $c_{n}$. In Ref. Bhattacharjee:2022lzy , the coefficient denoted by $a_{n}$ corresponds to our $c_{n}$, and it is also established that $c_{n}$ is associated with the dissipation coefficient. Meanwhile, we observe that $b_{n}$ remains identical for both the closed and open systems, as it is determined solely by the Hamiltonian.
In Ref. Parker:2018yvk , it is clearly established that the dynamical system is an infinite, chaotic, many-body system as $b_{n}=\alpha n+\gamma$, where $\alpha$ and $\gamma$ are parameters that depend on the specific model. In our case, since $b_{n}\propto n$, the early universe can be regarded as an ideal chaotic system.
VI.1.1 $b_{n}$ and $\mu_{2}$
In this work, we adopt the notation of Ref. Li:2024kfm to perform numerical simulations of the coefficients $b_{n}$ and $\mu_{2}$. We begin by examining the properties of $c_{n}$ within our framework. Recall that Eq. (9) (which is essentially the Schrödinger equation) involves only the Lanczos coefficients. The parameter $c_{n}$ will subsequently appear in the following modified equation:
@@ VI.1.1 $b_{n}$ and $\mu_{2}$
Here, we have defined $\tilde{c}_{n}=-ic_{n}$. In taking the continuous limit, we have employed the approximations $b_{n+1}\approx b_{n}$ and $\phi_{n+1}-\phi_{n-1}\approx 2\partial_{n}\phi$. The coefficients $b_{n}$ and $c_{n}$ are defined in Eq. (58) and Eq. (57), respectively. Their general expressions are given by
@@
where $\mu_{1}$, $\mu_{2}$, $\beta$ are determined by various models and $\mu_{2}$ plays a role of dissipative coefficient as shown in Ref. Li:2024kfm . In our work, one can explicitly find $\beta=1$ and combine Eqs. (57), (58), and (60), we find that the parameters $\mu_{1}$ and $\mu_{2}$ have the following forms,
@@
It should be noted that during inflation, the parameter $\mu_{2}$ is primarily determined by the wavenumber $k$ due to the slow-roll conditions. Upon entering the RD and MD eras, which is taking into account the effective mass, we observe that $\mu_{2}$ becomes smaller than unity, because the effective mass associated with these three types of potentials is much smaller than the momentum $k$.
We first discuss the Lanczos coefficient $b_{n}$. Figure 6 clearly illustrates the three distinct behaviors of $b_{n}$ during inflation, RD, and MD, where the parameters $k$ and $H_{0}$ are set to the same values as in the previous plots. As established in Ref. Bhattacharya:2022gbz , the Lyapunov exponent $\lambda$ is related to $b_{n}$ by $\lambda=2\alpha$, given the linear form $b_{n}=\alpha n+\delta$ (with $\delta=0$ in our case). This explicitly shows that $\lambda=2\alpha$, or equivalently $\lambda=b_{n}$ for $n=2$, as will be demonstrated in Figure 6. Physically, $\lambda$ characterizes the chaotic nature of the dynamical system. Consequently, in our framework, $b_{n}$ also reflects the chaotic behavior of the early universe.
@@
Figure 6 illustrates the evolution of the coefficient $b_{n}$ during inflation, RD, and MD. Based on the analysis of $b_{n}$, the degree of chaos increases due to the exponential expansion of the background spacetime. As the energy scale decreases, $b_{n}$ decreases correspondingly with the temperature. During MD, particles are produced via the preheating mechanism known as parametric resonance, leading to an increase in $b_{n}$.
Finally, we analyze the dissipation coefficient $\mu_{2}$, which is related to $c_{n}$, thereby confirming previous results from Ref. Li:2024kfm . Figure 7 clearly shows the evolution of $\mu_{2}$ in RD and MD. The value of $\mu_{2}$ during inflation is not shown in Fig. 7 because $\mu_{2}=k=1$ for the chosen wavenumber $k$ in the inflationary scenario. As asserted in Ref. Li:2024kfm , inflation constitutes a strong dissipative system with $\mu_{2}\geq 1$, while RD and MD are weak dissipative systems with $\mu_{2}\ll 1$, a result explicitly corroborated by Fig. 7.
@@
VI.2 General discussion of $\phi_{n}$
Before investigating the Krylov complexity, we discuss the general properties of $\phi_{n}$ since the definition of Krylov complexity is explicitly represented by $C_{K}=\sum_{n=1}^{+\infty}n|\phi_{n}|^{2}$. We will follow the general discussion of $\phi_{n}$ via Refs. Li:2024kfm ; Bhattacharjee:2022lzy . Eq. (59) is rewritten by
@@ VI.2 General discussion of $\phi_{n}$
where we have defined $\tilde{c}_{n}=n\chi\mu$ and $b_{n}=n\alpha$. As $\eta$ approaches infinity, the solution of Eq. (62) has the form $\phi(n)\propto e^{-n\xi}$ with $\xi=\frac{2\alpha}{\mu\chi}$, where this form will appear in RD or MD due to their weak dissipative nature. In the weak dissipative region, the corresponding Krylov complexity behaves as follows: it increases exponentially at the onset, with the average position given by $C_{k}\propto e^{2\alpha\eta}$ before reaching $\eta_{*}$. After this point, the Krylov complexity saturates to a constant value $\xi$ when $\eta>\eta_{*}$, where $\eta_{*}\sim\frac{1}{2\alpha}\ln{\frac{2\alpha}{\chi\mu}}$ dubbed as the scrambling time. It is evident that $\eta_{*}$ determines the range of exponential growth of Krylov complexity, determined by the dynamical system itself. It is straightforward to demonstrate that $\exp(2a\eta_{*})=\xi$. After some simple algebra, we obtain the formula $\xi$ in our case as follows,
@@
From this formula, we observe that $\xi$ is not strictly constant; it varies with respect to $a$ and $k$. Consequently, a significant change in the background will have a substantial impact on Krylov complexity. In this section, we examined the general properties of Krylov complexity without providing the precise formula for $\phi_{n}$. To fully capture the information concerning Krylov complexity within the framework of the open-system methodology, it is essential to have the exact wave function for $\phi_{n}$.
VI.3 Wave function within open-system methodology
Although we constructed the wave function for an open system in reference Li:2024kfm , based on the framework established in Bhattacharjee:2022lzy , this wave function cannot capture the complete information since it is the function of the conformal time $\eta$. To clearly illustrate this issue, we will first present our constructed wave function as follows:
@@ VI.3 Wave function within open-system methodology
To perform the Taylor expansion around $\mu_{2}$ to the leading order $(\mu_{2}=0)$, we obtain the expression
$\phi_{n}\approx\text{sech}(\eta)\,\tanh^{n}(\eta)+\mathcal{O}(\mu_{2}^{n}).$
To relate with Eq, (41), we can find one explicit correspondence between $r_{k}$ and $\eta$. To include the parameter of $\phi_{k}$, we will follow the guidelines provided in Appendix D of Parker:2018yvk to rigourisly derive the so-called open two-mode squeezed state, which will accurately reproduce the wave function represented in (41).
@@
where $x$ represents the Hamiltonian, $\tilde{c}_{n}$ is associated with the dissipative coefficient, and $b_{n}$ refers to the Lanczos coefficient. The sequence begins with the initial conditions $P_{0}(x)=1$ and $P_{1}(x)=x-c_{0}$. For an open system, we have $P_{n}(x;\alpha,\eta)=\det(x-\mathcal{L}_{n})$, where $\mathcal{L}_{n}$ is the Hamiltonian for the $n$-th quantum state. This determinant encodes information about $b_{n}$ as represented by the corresponding matrix. Based on the definition of $L$, we express it in terms of $x^{n}$ or $e_{n}$, which will be used to derive the wave function using the open-system methodology.
To obtain the similar recurrence relation as (53), we introduce the the natural orthonormal basis $\{e_{n}\}$, and meanwhile $P_{n}$ is represented in terms of $e_{n}$
@@
where $x$ is the parameter of $P_{n}$, $\mathcal{L}$ represents our Hamiltonian expressed in terms of creation and annihilation operators, and $|e_{n}\rangle\equiv|\mathcal{O}_{n}\rangle=\Bigl|n_{\vec{k}};n_{-\vec{k}}\Bigr\rangle$. By combining Equation (65) with Equation (66), we derive the following recurrence relation.
@@
where $\mu_{n,k}$ refers to a triangular linear transform with matrix elements that relate $x$ to $P_{n}$. This formula demonstrates that the polynomials $P_{n}$ is expressed in terms of $x$. All three orthogonal polynomial sequences (OPS) mentioned are examples of Sheffer orthogonal polynomials, which are defined as a set of polynomials $\{P_{n}(x)\}_{n\geq 0}$ characterized by a generating function,
@@
In our calculations, we multiply a phase factor $ie^{2i\phi_{k}(\eta)}$ which does not alter the wave function. Finally, we set $r_{k}=\eta$, allowing us to denote the two-mode squeezed state within the open system as,
@@
This wave function constitutes the central result of this paper. Its leading-order expansion of $\phi_{n}$ in terms of $\mu_{2}$ is consistent with the wave function presented in (41). The idea of performing a variable change is inspired by the construction of coherent states, where a key element is the displacement operator, defined as
$D(\eta)=e^{iHt},$
with $H$ being the corresponding Hamiltonian, and where the parameter $t$ can be re-expressed in terms of $r_{k}$ and $\phi_{k}$. As noted in Ref. Caputa:2021sib , a generalized coherent state can be derived by constructing a generalized displacement operator based on the appropriate group representation theory.
VI.4 The power of wave function (83)
We discuss the two-mode squeezed state in an open system, as described in equation (83). In contrast to the two-mode squeezed state derived from closed-system methedology, our wave function (83) is more physically realistic, since it includes the contribution of open system via Hamiltonian. Furthermore, although our derivation is obtained within the context of the early universe, this wave function is also widely applicable in fields such as quantum optics, quantum information, and condensed matter physics, where physical information is encoded in the parameters $\mu_{1}$ and $\mu_{2}$, which are determined by the specific system Hamiltonian. Finally, we verify the validity of our results by requiring that the recurrence relation (65), which is equivalent to the Lanczos algorithm, be satisfied. This connection underscores the utility of the Lanczos algorithm, which offers a systematic framework for constructing the wave function directly from the Hamiltonian, even for open quantum systems.
VI.5 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$ in open-system methodology
Following the standard procedure, we derive the evolution equations for $r_{k}$ and $\phi_{k}$ with respect to conformal time $\eta$. The details of this derivation are provided in Appendix A. The resulting evolution is summarized by the following equations: